# Direct integration of atomic precision advanced manufacturing into middle-of-line silicon fabrication

### v1

@@ III APAM sample preparation
Accommodating the limitations of the MOL of Sandia’s 0.35 CMOS will require lowering the APAM thermal budget to 600 ∘C. Fortunately, the only step of the typical APAM process that requires temperatures above 600 ∘C is removing the surface oxide through thermal desorption.[33, 34] The surface preparation needs to produce a surface that is sufficiently crystalline to dissociate phosphine, canonically requiring 3 neighboring dimers of the 21 reconstruction to be present.[35] A popular choice for low-temperature oxide removal is chemical etching using aqueous hydrofluoric acid, but this produces a surface contaminated with carbon. [36] While many alternatives exist, including vapor hydrofluoric acid etching,[36] we chose noble gas ion sputtering as a proxy for dry etch tools common in industry. Combined with an anneal to recrystallize the amorphized surface, sputtering has been used to produce sufficiently flat and clean surfaces for STM imaging.[37] The recrystallization process, called solid phase epitaxy (SPE), proceeds at rates of 1 nm/min to 10s of nm/min at temperatures between 500 ∘C and 600 ∘C for Si(100),[38] making it a good candidate given the available thermal budget. A similar annealing process is already used in APAM device fabrication - the locking layer process used to minimize dopant segregation during epitaxy (Section VII.2).[20, 19] For the locking layer process, a thin silicon layer is deposited after doping the surface at ambient temperature, trapping most of the dopant atoms within this layer by limiting adatom-mediated diffusion during silicon deposition. A rapid anneal improves the crystallinity of the layer while only activating the much slower process of bulk diffusion.
@@
In Figure 3(a), using high-angle annular dark field (HAADF) scanning transmission electron microscopy (STEM), we show the crystallinity of a sample sputtered for 30 min with neon ions accelerated at 2.0 keV at 60∘ from surface normal, and then annealed at 550-600 ∘C for 15 s to complete SPE. The phosphorus delta layer is discernible in the cross-sectional STEM image, [39] and since it is incorporated on the surface, it marks the depth of the surface after sputtering and annealing in the cross section. A buried layer of defects is visible approximately 5 nm below the phosphorus delta layer, possibly from damage at the end-of-range for 2 keV neon ions.[40] This buried layer marks the extent of the observed damage caused by our sputter process, meaning that a 5 nm layer of silicon was amorphized and then recrystallized. SPE of amorphous silicon proceeds at 20-30 nm/min at 575 ∘C,[38] which is consistent with this 15 s anneal recrystallizing 5 nm of silicon. Further optimization through use of lower energy ions can reduce the thickness of the amorphous layer, and the amount of annealing needed to repair it. This will likely bring the buried defect layer to the surface, which may prove desirable if that can be healed through SPE, or undesirable if it cannot.
@@
To establish a stronger correlation between crystallinity and electrical activation, we use STM to image the degree to which annealing heals the surface, and compare these data to magnetotransport measurements after delta layers are then made with these samples. These samples were sputtered with 1.5 keV neon ions, where the lower energy will produce roughly 75% of the amorphization depth in Figure 3(a). Annealing the sputtered sample for 1 min at 485 ∘C leaves the surface amorphous with no discernible atomic-scale ordering (Figure 3(b)), which is consistent with the low SPE rate of 0.6 nm/min.[38] Subsequently fabricated etch-defined Hall bars (Section VII.3) were not conductive at 4 K. Increasing the annealing time at 485 ∘C to 15 min produces a rough, but crystalline, surface where the dimer rows of the 21 reconstruction are visible on small terraces, despite significant topographical variation (Figure 3(c)). Qualitatively similar results were obtained for 5 min at 500 ∘C, which is consistent with the higher SPE rate of 0.9 nm/min at this temperature.[38] Etch-defined Hall bars made from these rough but crystalline surfaces all produce a carrier density in excess of 1.2 cm-2 at 4 K. These results demonstrate that a high carrier density is correlated with the creation of a crystalline surface.
@@
Going further to 560 ∘C and 600 ∘C for 15 minutes (Figure 3(d)), produced a crystalline surface with larger terraces of dimer rows than lower temperatures, but also revealed surface defects (e.g. in the upper-right corner of the image) that might arise from contaminants in the buried defect layer diffusing to the surface. For samples annealed for 15 min at different temperatures, the 2D carrier density plateaus after exceeding the threshold annealing temperature where the surface becomes crystalline (Figure 4(a)). A surface where the dimer rows of the Si(100) 21 reconstruction are present is likely all that is needed for the decomposition pathways to proceed as proposed for various dopant precursors.[35, 41, 42] Conversely, an amorphous surface does not contain the ordered dimers used in these decomposition pathways, reducing dopant incorporation below the threshold for metallic conduction, 3.7 cm-3 at 4 K.[43] The mobility continues to improve from 22 cm2V-1s-1 to 32 cm2V-1s,-1 and thus the resistance continues to decrease, on increasing the annealing temperature from 485 ∘C to 600 ∘C (Figure 4(b)). This improvement in mobility correlates with a decrease in the root mean square (RMS) surface roughness of the layer from 274 pm to 167 pm, with the caveat that the smoothest layers examined here, ranging 123-132 pm, are amorphous and thus insulating at 4 K. To what extent this trend in mobility is purely a function of the roughness or is also influenced by the buried defect layer will require further work to determine. For comparison, our typical APAM process with a high-temperature surface clean produces material with similar densities (1.5-2.0 cm-2) and even higher mobilities (40-50 cm2V-1s-1), consistent with published results.[19] That material has an even lower RMS surface roughness (50 pm or less), and has no buried layer of defects from the sputter process.
@@ IV APAM lithography
The magnetotransport data from a Hall bar defined using a thicker 1 composite dielectric, similar to our CMOS chips, shows both visible and electrical signatures of shadowing. An optical image, Figure 6(a), shows discoloration extending 3 from the hardmask where oxide was never removed, and dopants never incorporated. The size of this shadowed region is nearly double that expected from the hardmask thickness of 1 and a sputter angle of 60∘, but is also clearly visible from the optical image. The magnetotransport data was first analyzed using the assumed geometry, where this shadowing is not accounted for, and is shown as red triangles in Figure 6(b-c). Shadowing was seen to shrink the width of the Hall bar from 20 to 17 , and increase the number of squares by nearly 20%. The thick hardmask sample has a carrier density similar to the thin hardmask samples, both less than the samples from Section III. Once more, this might arise either from redeposition from the hardmask wall during sputtering, or the use of a dirtier PECVD oxide instead of a cleaner thermal oxide. Conversely, these thick hardmask samples have similar mobility to the samples in Section III, both lower than the thin hardmask sample. This might come from both being prepared using 1.5 keV neon ions, and having a shallower defect layer than produced by 2.0 keV ions in the case of the thin hardmask sample. A shallow defect layer might scatter carriers more, reducing mobility.
@@ V CMOS + APAM
Finally, we characterize the NMOS with the APAM resistor connected in series with the drain using the APAM in Series with Drain bond pad (red dot). The turn-on behavior in Figure 8(c), red line, remains consistent for both with the only change being a flattening of the drain current at a gate bias of 1.5 V with the APAM resistor. This maximum drain current indicates when the resistance of the APAM resistor exceeds the channel resistance. We can confirm this by taking the slope of the output characteristics at a gate bias of 3 V (Figure 8(d), red line) which gives a resistance of approximately 13 k. The majority of the 13 k can be accounted for by the APAM resistor with the rest coming from the NMOS and associated parasitics. At a lower gate bias of 1.8 V (blue line), the drain current goes from being limited by the APAM resistor to being limited by saturation in the channel above 2 V. Finally, at an even lower gate bias of 1.2 V, the drain current is limited by saturation in the channel for much of the range of drain bias. This successful integration and understanding highlights the compatibility of our modified APAM process with CMOS to exploit CMOS + APAM for future microelectronic applications.
@@ VI Conclusion
The significance of this proof-of-principle integration of APAM and CMOS is both conceptual and practical. The first advance is the articulation of a detailed methodology for direct integration of a boutique process. This is often too risky when existing microfabrication tools cannot recreate the desired process in the manufacturing facility. The inability to access the details of proprietary CMOS workflows can also make ambitious direct integration inaccessible to researchers. The roadmap used here can likely be extended to other boutique processes which have similar thermal envelopes, and are being integrated with other manufacturing workflows, including modern ones. The second advance is the specification of a tool capable of bringing APAM out of the laboratory and into the microfabrication facility, expanding access from a limited number of domain experts to a much wider community. In practice, we anticipate that a more optimized process will be needed by industry given access to technologies which are common in manufacturing but are inaccessible at the laboratory scale, including deep and extreme ultraviolet photolithography, proprietary low-temperature dielectric etches, ultrafast rapid thermal annealing, and low-temperature chemical vapor deposition. Conversely, an integrated tool which can perform a sputter and anneal, chemical doping, and silicon homoepitaxy is simple and flexible enough to incorporate into university-scale clean rooms to spur innovation. Finally, the potential impact of APAM-CMOS integration can best be imagined by looking at the design space for semiconductor electronic structure enabled by alloying silicon with germanium, and the large number of devices this has inspired. We envision that the wider availability of APAM will result in the device community no longer thinking of doping as just adding either electron or hole carriers to the parent semiconducting material, but doped material having novel properties itself. Only then can the design space for doped material serve as a source of inspiration for the discovery of new devices.
@@ VII.2 APAM processing
Samples undergoing typical high-temperature APAM processing were degassed at 550 ∘C for 20 min with a pBN heating element to make the silicon sufficiently conductive to easily carry a DC current (a standard APAM heating process) and then at 750 ∘C for 40 min via DC Joule heating. Thermal desorption of the 3 nm of oxide was accomplished through 6 cycles alternating between 750 ∘C and 950 ∘C-1050 ∘C for a total of 60 min at 950 ∘C-1050 ∘C by DC Joule heating, with temperatures measured by optical pyrometry.

### v2

@@ III APAM sample preparation
Accommodating the limitations of the MOL of Sandia’s 0.35 CMOS will require lowering the APAM thermal budget to 600 ∘C. Fortunately, the only step of the typical APAM process that requires temperatures above 600 ∘C is removing the surface oxide through thermal desorption.[33, 34] The surface preparation needs to produce a surface that is sufficiently crystalline to dissociate phosphine, canonically requiring 3 neighboring dimers of the 21 reconstruction to be present.[35] A popular choice for low-temperature oxide removal is chemical etching using aqueous hydrofluoric acid, but this produces a surface contaminated with carbon. [36] While many alternatives exist, including vapor hydrofluoric acid etching,[36] we chose noble gas ion sputtering as a proxy for dry etch tools common in industry. Combined with an anneal to recrystallize the amorphized surface, sputtering has been used to produce sufficiently flat and clean surfaces for STM imaging.[37] The recrystallization process, called solid phase epitaxy (SPE), proceeds at rates of 1 nm/min to 10s of nm/min at temperatures between 500 ∘C and 600 ∘C for Si(100),[38] making it a good candidate given the available thermal budget. In SPE, silicon atoms at the amorphous-crystal boundary break bonds, diffuse a short distance, and remake bonds into a more ideal crystalline configuration. [38] This atomic rearrangement happens at a lower temperature than dopant diffusion in the bulk, which is driven by interstitial diffusion and then vacancy diffusion above 600 ∘C, [39, 40] and might be an important factor in maintaining transistor electrical characteristics. A similar annealing process is already used in APAM device fabrication - the locking layer process used to minimize dopant segregation during epitaxy (Section VII.2).[20, 19] For the locking layer process, a thin silicon layer is deposited after doping the surface at ambient temperature, trapping most of the dopant atoms within this layer by limiting adatom-mediated diffusion during silicon deposition. A rapid anneal improves the crystallinity of the layer. Dopant diffusion is more likely to occur in this thin silicon layer during the anneal than in the bulk due to the motion of a high density of vacancies and interstitials formed during the low temperature epitaxy. However, as shown in Section II, after crystallization and additional growth, the thermal budget of the buried APAM layer resembles dopants in the bulk.
@@
In Figure 3, using high-angle annular dark field (HAADF) scanning transmission electron microscopy (STEM), we show the crystallinity of a sample sputtered for 30 min with neon ions accelerated at 2.0 keV at 60∘ from surface normal, and then annealed at 550-600 ∘C for 15 s to complete SPE. The phosphorus delta layer is discernible in the cross-sectional STEM image, [41] and since it is incorporated on the surface, it marks the depth of the surface after sputtering and annealing in the cross section. A buried layer of defects is visible approximately 5 nm below the phosphorus delta layer, possibly from damage at the end-of-range for 2 keV neon ions.[42] This buried layer marks the extent of the observed damage caused by our sputter process, meaning that a 5 nm layer of silicon was amorphized and then recrystallized. SPE of amorphous silicon proceeds at 20-30 nm/min at 575 ∘C,[38] which is consistent with this 15 s anneal recrystallizing 5 nm of silicon. Further optimization through use of lower energy ions can reduce the thickness of the amorphous layer, and the amount of annealing needed to repair it. This will likely bring the buried defect layer to the surface, which may prove desirable if that can be healed through SPE, or undesirable if it cannot.
@@
To establish a stronger correlation between crystallinity and electrical activation, we use STM to image the degree to which annealing heals the surface, and compare these data to magnetotransport measurements after delta layers are then made with these samples. These samples were sputtered with 1.5 keV neon ions, where the lower energy will produce roughly 75% of the amorphization depth in Figure 3. Annealing the sputtered sample for 1 min at 485 ∘C leaves the surface amorphous with no discernible atomic-scale ordering (Figure 4(a)), which is consistent with the low SPE rate of 0.6 nm/min.[38] Subsequently fabricated etch-defined Hall bars (Section VII.3) were not conductive at 4 K. Increasing the annealing time at 485 ∘C to 15 min produces a rough, but crystalline, surface where the dimer rows of the 21 reconstruction are visible on small terraces, despite significant topographical variation (Figure 4(b)). Qualitatively similar results were obtained for 5 min at 500 ∘C, which is consistent with the higher SPE rate of 0.9 nm/min at this temperature.[38] Etch-defined Hall bars made from these rough but crystalline surfaces all produce a carrier density in excess of 1.2 cm-2 at 4 K. These results demonstrate that a high carrier density is correlated with the creation of a crystalline surface.
@@
Going further to 560 ∘C and 600 ∘C for 15 minutes (Figure 4(c)), produced a crystalline surface with larger terraces of dimer rows than lower temperatures, but also revealed surface defects (e.g. in the upper-right corner of the image) that might arise from contaminants in the buried defect layer diffusing to the surface. For samples annealed for 15 min at different temperatures, the 2D carrier density plateaus after exceeding the threshold annealing temperature where the surface becomes crystalline (Figure 5(a)). A surface where the dimer rows of the Si(100) 21 reconstruction are present is likely all that is needed for the decomposition pathways to proceed as proposed for various dopant precursors.[35, 43, 44] Conversely, an amorphous surface does not contain the ordered dimers used in these decomposition pathways, reducing dopant incorporation below the threshold for metallic conduction, 3.7 cm-3 at 4 K.[45] The mobility continues to improve from 22 cm2V-1s-1 to 32 cm2V-1s,-1 and thus the resistance continues to decrease, on increasing the annealing temperature from 485 ∘C to 600 ∘C (Figure 5(b)). This improvement in mobility correlates with a decrease in the root mean square (RMS) surface roughness of the layer from 274 pm to 167 pm, with the caveat that the smoothest layers examined here, ranging 123-132 pm, are amorphous and thus insulating at 4 K. To what extent this trend in mobility is purely a function of the roughness or is also influenced by the buried defect layer will require further work to determine. For comparison, our typical APAM process with a high-temperature surface clean produces material with similar densities (1.5-2.0 cm-2) and even higher mobilities (40-50 cm2V-1s-1), consistent with published results.[19] That material has an even lower RMS surface roughness (50 pm or less), and has no buried layer of defects from the sputter process.
@@ IV APAM lithography
The magnetotransport data from a Hall bar defined using a thicker 1 composite dielectric, similar to our CMOS chips, shows both visible and electrical signatures of shadowing. An optical image, Figure 7(a), shows discoloration extending 3 from the hardmask where oxide was never removed, and dopants never incorporated. The size of this shadowed region is nearly double that expected from the hardmask thickness of 1 and a sputter angle of 60∘, but is also clearly visible from the optical image. The magnetotransport data was first analyzed using the assumed geometry, where this shadowing is not accounted for(uncorrected width in Figure 7(a)), and is shown as red triangles in Figure 7(b-c). Shadowing was seen to shrink the width of the Hall bar from 20 to 17 , and increase the number of squares by nearly 20%. Analyzing the magnetotransport data using this corrected width (Figure 7(a)) yielded the green squares in Figure 7(b-c), with negligible impact on the electron density, but a 25% increase in mobility compared to the uncorrected width. The thick hardmask sample has a carrier density similar to the thin hardmask samples, both less than the samples from Section III. Once more, this might arise either from redeposition from the hardmask wall during sputtering, or the use of a dirtier PECVD oxide instead of a cleaner thermal oxide. Conversely, these thick hardmask samples have similar mobility to the samples in Section III, both lower than the thin hardmask sample. This might come from both being prepared using 1.5 keV neon ions, and having a shallower defect layer than produced by 2.0 keV ions in the case of the thin hardmask sample. A shallow defect layer might scatter carriers more, reducing mobility.
@@ V CMOS + APAM
Finally, we characterize the NMOS with the APAM resistor connected in series with the drain using the APAM in Series with Drain bond pad (red dot). The turn-on behavior in Figure 9(c), red line, remains consistent for both with the only change being a flattening of the drain current at a gate bias of 1.5 V with the APAM resistor. This maximum drain current indicates when the resistance of the APAM resistor exceeds the channel resistance. We can confirm this by taking the slope of the output characteristics at a gate bias of 3 V (Figure 9(d), red line) which gives a resistance of approximately 13 k. The majority of the 13 k can be accounted for by the APAM resistor with the rest coming from the NMOS and associated parasitics. At a lower gate bias of 1.8 V (blue line), the drain current goes from being limited by the APAM resistor to being limited by saturation in the channel above 2 V. Finally, at an even lower gate bias of 1.2 V, the drain current is limited by saturation in the channel for much of the range of drain bias. We acknowledge that the proof-of-concept demonstrated here had low yield of functional CMOS + APAM devices, primarily due the shadow mask discussed above. We anticipate that adapting this hardmask approach to industry-standard tools for lithography, etch, epitaxy, and ohmic contacts would ultimately reach device yields consistent with that of the foundry implementing the APAM doping after sufficient process development. This successful integration and understanding highlights the compatibility of our modified APAM process with CMOS to exploit CMOS + APAM for future microelectronic applications.
@@ VI Conclusion
The significance of this proof-of-principle integration of APAM and CMOS is both conceptual and practical. The first advance is the articulation of a detailed methodology for direct integration of a boutique process. This is often too risky when existing microfabrication tools cannot recreate the desired process in the manufacturing facility. The inability to access the details of proprietary CMOS workflows can also make ambitious direct integration inaccessible to researchers. The roadmap used here can likely be extended to other boutique processes which have similar thermal envelopes, and are being integrated with other manufacturing workflows, including modern ones. The second advance is the specification of a tool capable of bringing APAM out of the laboratory and into the microfabrication facility, expanding access from a limited number of domain experts to a much wider community. In practice, we anticipate that a more optimized process will be needed by industry given access to technologies which are common in manufacturing but are inaccessible at the laboratory scale, including deep and extreme ultraviolet photolithography, proprietary low-temperature dielectric etches, ultrafast rapid thermal annealing, and low-temperature chemical vapor deposition. The high throughput a manufacturing process enables is a pre-requisite for yield exploration or optimization, and is unlikely to be related to the bespoke process detailed here. Conversely, an integrated tool which can perform a sputter and anneal, chemical doping, and silicon homoepitaxy is simple and flexible enough to incorporate into university-scale clean rooms to spur innovation. Finally, the potential impact of APAM-CMOS integration can best be imagined by looking at the design space for semiconductor electronic structure enabled by alloying silicon with germanium, and the large number of devices this has inspired. We envision that the wider availability of APAM will result in the device community no longer thinking of doping as just adding either electron or hole carriers to the parent semiconducting material, but doped material having novel properties itself. Only then can the design space for doped material serve as a source of inspiration for the discovery of new devices.
@@ VII.2 APAM processing
Samples undergoing typical high-temperature APAM processing were degassed at 550 ∘C for 20 min with a pBN heating element to make the silicon sufficiently conductive to easily carry a DC current (a standard APAM heating process) and then at 750 ∘C for 40 min via DC Joule heating. Thermal desorption of the 3 nm of oxide was accomplished through 6 cycles alternating between 750 ∘C and 950 ∘C-1050 ∘C for a total of 60 min at 950 ∘C-1050 ∘C by DC Joule heating, with temperatures measured by optical pyrometry. While higher temperatures are ordinarily used to ensure that no carbon remains on the surface, there was no evidence of carbon contamination during subsequent STM imaging. This results from assuring photoresist is only baked in contact with thermally oxidized silicon during microfabrication, and using an ex situ oxygen plasma treatment before thermal oxide desorption.